# Measurement-induced asymmetry in bipartite networks

### v1

@@ Abstract
We consider an interacting bipartite network through a Bayesian game-theoretic framework and demonstrate that weak measurements introduce an inherent asymmetry that is not present when using standard projective measurements. These asymmetries are revealed in the expected payoff for a Bayesian version of Prisoner’s dilemma, demonstrating that certain advantages can be found for given subsystems depending on the measurements performed in the network. We demonstrate that this setup allows measurement-induced control for one of the respective parties.
@@ I Introduction
Measurement theory is an increasingly active area of research, from metrology [1, 2, 3, 4], to understanding the fundamental nature of systems [5, 6]. However, when considering quantum systems, the standard approach of using projective measurements only reveals certain aspects of the model due to the huge disturbance they produce (the so-called back-action). To overcome this problem, weak measurements [7] are used as their back-action is small enough to perturb the measured system barely. As a trade-off, little information about the system is gained. More precisely, performing a weak measurement aims to extract a partial amount of information from a state without entirely reducing it to an eigenstate of a given observable. Weak measurements have been implemented to prepare quantum states [8, 9, 10, 11], improve the efficiency in state tomography [12, 13], simultaneously measure non-commuting observables [14, 15, 16], reveal more correlation in multipartite systems [17], and to amplify of signals in quantum metrology [18, 19, 20] utilizing the closely related concept of weak-values [21], which generalize the concept of expectation values in quantum mechanics, among other applications [22].
@@
Alongside this, the understanding of interacting parties in a network [23, 24, 25, 26] is relevant to broader society due to the increasing number of emerging interacting networks. A host of different techniques are used to understand the underlying networks, depending on the network’s topology [27, 28]. Understanding how each party can affect the overall system with their local measurements and operations is vital to understanding the overall nature of the system and the correlation between the parties [29, 30].
@@
A natural framework to understand and model how parties can interact is through game-theoretic techniques [31, 32]. This results in a scenario where the two parties perform certain operations and measurements in order to optimize their respective situation. These types of scenarios occur in everyday life, from buying and selling goods to determining the best route of a journey [33, 34, 35].
@@
It is well established that asymmetry can be present in correlations between interacting parties, either through quantum discord [36, 37], or standard classical correlation. Asymmetry in this context relates to the asymmetric correlation between two parties. For example, the quantum discord of a given state can differ depending on which subsystem is measured and optimized over. Despite this non-trivial aspect, understanding the link between measurement theory and asymmetry has received little attention [38, 39]. Subsequently, the primary motivation of this work is to establish whether weak measurements allow any form of asymmetry to be introduced in an interacting network.
II Generalized Measurments
In order to define the measurements used throughout this work, it is important to introduce the physical motivation for how different measurements can be implemented and verified. This requires formally introducing weak measurements, and how they differ from projective measurements.
@@ II.1 Weak Measurements
    An ancillary system $D$ initially prepared in the state |$\ket{\phi}\in\mathcal{H}_{D}$ is coupled with $S$ via an interaction Hamiltonian
    $H_{\text{DS}}=\sum_{n}\sqrt{\frac{\gamma_{n}}{\delta t}}A_{n}\otimes B_{n},$ (1)
    where $\delta t$ is the interaction time, and $\gamma_{n}$ are measurement strengths with units of inverse time such that the weak-measurement condition $\sqrt{\gamma_{n}\delta t}=o(1)$ holds for all $n$. Furthermore, one can also assume that $J_{D}/(\gamma_{n}\delta t)=J_{S}/(\gamma_{n}\delta t)=o(1)$ is satisfied for all $n$ so the self-evolution of both the system and detector can be ignored. Here, $J_{D}$ and $J_{S}$ are the energy scales of the detector and system Hamiltonians, respectively [40].
@@
    The total system $S+D$, which initially is a product state |$\ket{\phi}\otimes\ket{\psi}$ evolves during a time $\delta t$ and the ancilla is projectively measured on a given basis {|$\{\ket{\alpha}\}$. Therefore, the updated state of $S$ is given by
@@
    where the set of Kraus operators $\{M_{\alpha}\}=\{\bra{\alpha}\exp(-i\delta tH_{\text{DS}})\ket{\phi}\}$ describes every possible indirect measurements on $S$ given the interaction (1), the state |$\ket{\phi}$ and the measurement basis {|$\{\ket{\alpha}\}$. The Kraus operators satisfy $\sum_{\alpha}M_{\alpha}^{\dagger}M_{\alpha}=I_{S}$.
@@
    A post-selection of the states is performed. This step is required depending on the chosen measurement basis, as the back-action of the measured ancilla can induce an effective discontinuous evolution in (2); that is, under the weak-measurement condition, upon expanding $M_{\alpha}$ as a series in $\delta t$, depending on the post-selected state |$\ket{\alpha}\otimes\ket{\psi_{\alpha}}$, the latter does not evolve continuously from the initial state |$\ket{\phi}\otimes\ket{\psi}$. Therefore, the term weak comes from the (possibly) post-selected trajectories of the effectively evolved state that evolved continuously and non-unitarily from the prior state |$\ket{\psi}$.
@@ II.2 Meausurement operators
Let $\sigma,\sigma^{\prime}\in\{1,-1\},$ $\bm{\alpha}\in\{\bm{a},\bm{a}^{\prime}\}$ and $\bm{\beta}\in\{\bm{b},\bm{b}^{\prime}\},$ where $\bm{a},\bm{a}^{\prime},\bm{b},\bm{b}^{\prime}\in\mathcal{S}^{2}$ are fixed Bloch vectors parameterized in spherical coordinates. We define the generalized measurements acting on 4-by-4 density matrices via the following Kraus (or measurement) operators [41, 42]
@@
where $P(\bm{\alpha},\bm{\beta})$ is some joint probability density of obtaining $(\bm{\alpha},\bm{\beta})$ and
$M_{\pm 1|\bm{\alpha}}(y)=a(\pm y)\Pi_{+|\bm{\alpha}}+a(\mp y)\Pi_{-|\bm{\alpha}},$ (4)
are also Kraus operators with
@@
as the projector associated with the Bloch vector $\bm{\alpha}=(\sin\theta_{\alpha}\cos\phi_{\alpha},\sin\theta_{\alpha}\sin\phi_{\alpha},\cos\theta_{\alpha})$, and $\bm{\sigma}$ is the vector of Pauli matrices. Furthermore,
@@
where $y$ characterizes the strength of the measurement. For example, for small $y$, the measurement is denoted as weak. In particular, if $y=0$, then $M_{\sigma|\bm{\alpha}}(0)=I_{2}/\sqrt{2}$, where $I_{2}$ denotes the 2-by-2 identity, and the measurement does nothing to the state $\rho$. In contrast, if $y\rightarrow\infty$, (4) becomes a single projector; thus, this is the strong-measurement limit. All the above definitions hold as well for $\sigma^{\prime}$ and $\bm{\beta}$.
Finally, it is straightforward to verify that Kraus operators in (3) are well-defined for any $y$ and $z$. First, they resolve the identity in $\mathbb{C}^{4}$
@@
for an arbitrary density matrix $\rho$. Furthermore, since we defined (6) as a positive linear combination of projectors, it implies that the Kraus operators are completely positive [42], i.e., $I_{4}\otimes K^{\dagger}(\bm{\alpha},\bm{\beta};\sigma,\sigma^{\prime};y,z)K(\bm{\alpha},\bm{\beta};\sigma,\sigma^{\prime};y,z)>0$.
@@ III Bayesian Game Theory
The game-theoretic focus of this work is on Bayesian or incomplete information games. These were formulated classically by Harsanyi [43] and have been extended in recent years to include quantum mechanics [44, 45, 46, 47]. The essence is that players may be playing various games but are uncertain (i.e., they can have differing priors) about which game they are playing. Therefore, the probability distributions between players in a Bayesian game are conditioned upon which game they are playing and their subsequent choices. The conditional probability and priors determine the expected payoff for the players, where the actual payoffs can be arbitrary real numbers but are predetermined and known to both players. Importantly, there is a preference relation for these payoffs for the players. Throughout this paper, the largest number is preferred for the players; therefore, three is preferred over two. Subsequently, each player wishes to maximize their respective expected payoff.
In this work, we will consider an incomplete game of two players $A$ and $B$ sharing a quantum state $\rho$ describing a two-qubit system, where each player can perform local measurements represented by (4) on their subsystem. The parameters associated with player $A$ and $B$ are $\{\sigma,\bm{\alpha},y\}$ and $\{\sigma^{\prime},\bm{\beta},z\}$, respectively, where $\bm{a},\bm{a}^{\prime},\bm{b},\bm{b}^{\prime}$ and the strengths of the measurements $y$ and $z$ are fixed throughout the game. The incompleteness of this game is reflected in the fact that the players do not know beforehand in which direction they will be told to measure, e.g., in one trial, player $A$ can be told to measure $\bm{\alpha}=\bm{a}$ whereas player $B$ measures $\bm{\beta}=\bm{b}^{\prime}$. It is an independent referee who tells each player what direction to measure. After the measurements have been performed, the results of the measurements are communicated back to the referee, and respective payoffs are assigned. An expected payoff can be calculated once this procedure has been repeated many times, an expected payoff can be calculated.
@@
where |$\ket{x}=\cos(x/2)\ket{0}+\sin(x/2)\ket{1}$ and |$\ket{0},\ket{1}$ are in the computational basis. Given the above ingredients, the expected payoff [43] in our incomplete information game is given by
@@
where $u_{A,\sigma,\sigma^{\prime}}^{\bm{\alpha},\bm{\beta}}$ denotes a tensor of payoffs for Player $A$, $P(\bm{\alpha},\bm{\beta})$ denotes the prior probability, and the $\theta$’s denote the angles parameterizing the Bloch vectors. A similar expression holds for Player $B$.
@@ III.1 Measurement-induced asymmetry
In order to clarify what asymmetry means in this context, it is instructive to understand that by using projective measurements, the expected payoff in Bayesian game theory will depend on either players’ measurements or neither. Therefore, choosing a local measurement that renders Player $A$’s expected payoff independent of Player $B$’s measurement while ensuring that Player $B$’s expected payoff still depends on Player $B$’s measurements is unexpected. This difference in measurement dependence on the expected payoffs is considered asymmetry moving forward. This applies if players $A$ and $B$ are interchanged.
Collapsing the expected payoff such that it is independent of the other player’s measurements could have practical benefits for targeting a given value. This can be illustrated by using the Prisoner’s dilemma. The standard version of the game considers two prisoners who have been arrested for two offenses; one is a major offense (i.e., bank robbery), and the other is a minor offense (i.e., petty theft), but there is only enough evidence to charge them for the minor offense. However, the police would like to charge them for the major offense as well, but there is insufficient evidence of this. Therefore, the police gives the prisoners a choice, where each prisoner can give evidence incriminating the other prisoner in exchange for a better prison sentence. However, if they both incriminate each other, then each prisoner will serve a longer prison sentence. Crucially, neither prisoner knows what the other will do.
In what follows, we will consider a Bayesian version of the Prisoner’s dilemma with the added caveat that the prisoners do not know whether their respective police officer interviewing them is either honest (H) or corrupt (C).
Table 1: Bayesian version of the Prisoner’s dilemma. Q represents quiet, and T represents tell. The police officer’s motives and the prisoner’s choices can be mapped onto ${\bm{a}},{\bm{a}^{\prime}},{\bm{b}},{\bm{b}^{\prime}},+1$ and $-1$ directly in the measurement operators (4), e.g., for Prisoner A, Honest $\mapsto{\bm{a}}$, Corrupt $\mapsto{\bm{a^{\prime}}}$, Q $\mapsto+1$, and T $\mapsto-1$. For Prisoner B, Honest $\mapsto{\bm{b}}$, Corrupt $\mapsto{\bm{b}^{\prime}}$, Q $\mapsto+1$, and T $\mapsto-1$. Here, Prisoner A chooses the row choices, and Prisoner B chooses the column choices.
Since the corrupt police officer does not want to get caught, the officer rewards the prisoners for confessing and taking responsibility for the crime, i.e., tell (T), as this means that C will not get caught. The reward can be thought of as a financial payment, but the payoffs generally are arbitrary. However, if the Prisoners remain quiet (Q), then this increases the chance of C being caught, so the Prisoners do not get rewarded. A version of this type of game can be seen in Table 1.
@@
It can be seen that there is an asymmetry in the expected payoffs that arises due to the measurement when considering only one of the player’s payoffs. For example, considering Prisoner $A$’s payoff, Prisoner A could set $y=0$, which would collapse all the terms except $3/2$ in (12). Prisoner B does not have this level of control over Prisoner $A$’s payoff, as if $z=0$, there remain other terms that Prisoner $A$ controls. Interestingly, this asymmetry in the expected payoffs is reversed when considering the other player’s payoff in (13). Namely, Prisoner $B$ does not have the same level of control over Prisoner $A$’s payoff as Prisoner $A$.
It is clear that in the projective limit, the payoffs are symmetric in the sense that $u_{A}$ becomes $u_{B}$ when changing player A’s measurements to player B’s measurements, such that $\theta_{a}\rightarrow\theta_{b}$ and so on. This is not necessarily the case when weak measurements are included. Despite the asymmetry in the expected payoffs due to weak measurements, there is still the possibility of a given Prisoner optimizing their own respective payoff for a given choice of measurement. For example, Prisoner $A$ can set $y=0$ and fix the expected payoff at $3/2$. Despite this, Prisoner $B$ would still be left to control the measurements for the payoff $u_{B}$. Thus, in theory, Prisoner B would be able to acquire a larger expected payoff than Prisoner $A$ by choosing the detector settings ($\theta_{b},\theta_{b^{\prime}}$) in an optimal way and vice versa for Prisoner $B$. Subsequently, each player (if they so choose) has complete control over their respective expected payoff and can make their expected payoff completely independent of the other player’s choices (measurements) while allowing the other player to still have control over the corresponding particular payoff. This could have an application where one player is more risk-averse compared to the other, so this player can choose whether to allow the other player to influence their respective payoff.

### v2

@@ Abstract
We consider an interacting bipartite network through a Bayesian game-theoretic framework and demonstrate that weak measurements introduce an inherent asymmetry that is not present when using standard projective measurements. These asymmetries are revealed in the expected payoff for a Bayesian version of the Prisoner’s dilemma, demonstrating that certain advantages can be found for given subsystems depending on the measurements performed in the network. We demonstrate that this setup allows measurement-induced control for one of the respective parties.
@@ I Introduction
Measurement theory is an increasingly active area of research, from metrology [1, 2, 3, 4], to understanding the fundamental nature of systems [5, 6]. However, when considering quantum systems, the standard approach of using projective measurements only reveals certain aspects of the model due to the huge disturbance (or back action) they produce. To overcome this problem, weak measurements [7] are used as their back action can be small enough to perturb the measured system. As a trade-off, little information about the system is gained. More precisely, performing a weak measurement aims to extract a partial amount of information from a state without entirely reducing it to an eigenstate of a given observable. Weak measurements have been implemented to prepare quantum states [8, 9, 10], improve the efficiency in state tomography [11, 12], simultaneously measure non-commuting observables [13, 14, 15], reveal more correlation in multipartite systems [16], and to amplify signals in quantum metrology [17, 18, 19] utilizing the closely related concept of weak-values [20]. These generalize the concept of expectation values in quantum mechanics, among other applications [21].
@@
Alongside this, the understanding of interacting parties in a network [22, 23, 24, 25] is highly relevant across multiple scientific fields due to the increasing number of emerging interacting neworks. A host of different techniques are used to understand the underlying networks depending on its topology [26, 27]. Understanding how each party can affect the overall system with their local measurements and operations is vital to understanding the overall nature of the system and the correlation between the parties [28, 29].
@@
A natural framework to understand and model how parties can interact is through game-theoretic techniques [30, 31]. This results in a scenario where the two parties perform certain operations and measurements in order to optimize their respective situation. These types of scenarios occur in everyday life, from buying and selling goods to determining the best route of a journey [32, 33, 34], among many others.
@@
It is well established that asymmetry can be present in correlations between interacting parties, either through, e.g., quantum discord [35, 36], or standard classical correlation. Asymmetry in this context relates to the asymmetric correlation between two parties. For example, the quantum discord of a given state can differ depending on which subsystem is measured and optimized. Despite this non-trivial aspect, understanding the link between measurement theory and asymmetry has received little attention [37, 38]. Subsequently, the primary motivation of this work is to establish whether weak measurements induce any form of asymmetry to be introduced in an interacting network.
Our main contribution to the field is furthering the understanding of how measurement theory can be used as a mechanism of control by inducing asymmetry between two parties that have access to a shared state. In particular, this has been studied using game-theoretic techniques, which map to experimentally accessible frameworks [39, 40]. The proposed techniques could be used to design future interacting networks.
The paper is organized as follows. In Sec. II, we briefly describe the concept of generalized measurements and define the measurement operators we use in this work. In Sec. III, we introduce the notions of Bayesian games in the context of quantum mechanics and describe which type of quantum state is used. In Sec. III.1, we demonstrate a particular example of a Prisoner’s dilemma, i.e., we provide a given table of payoffs and allowed measurements, that lead to asymmetry in the expected payoffs for each prisoner when the prisoners implement a particular type of generalized measurement on their share of the quantum system. We also show that one prisoner can tune their measurements to ensure one of the prisoner’s expected payoffs is independent of the other prisoner’s detector settings.
II Generalized Measurements
In this section, we provide a brief summary of the concept of weak measurements, and define the measurement operators we use throughout this paper, which, under certain limits, correspond to either weak or projective measurements.
@@ II.1 Weak Measurements
    An ancillary system $D$ initially prepared in the state |$\ket{\phi}\in\mathcal{H}_{D}$ is coupled to $S$ via an interaction Hamiltonian
    $H_{DS}=\sum_{n}\sqrt{\frac{\gamma_{n}}{\delta t}}A_{n}\otimes B_{n},$ (1)
    where $\delta t$ is the interaction time, and $\{\gamma_{n}\}$ are measurement strengths with units of inverse time such that the weak-measurement condition $\sqrt{\gamma_{n}\delta t}=o(1)$ holds for all $n$. Furthermore, one can also assume that the conditions $J_{D}/(\gamma_{n}\delta t)=J_{S}/(\gamma_{n}\delta t)=o(1)$ are satisfied for all $n$ so the self-evolution of both the system and detector can be ignored. Here, $J_{D}$ and $J_{S}$ are the energy scales of the detector and system Hamiltonians, respectively [41].
@@
    The total system $S+D$, which initially prepared in the state |$\ket{\phi}\otimes\ket{\psi}$, evolves during a time $\delta t$ and the ancilla is projectively measured on a given basis {|$\{\ket{\alpha}\}$. Therefore, the updated state of $S$ is given by
@@
    where the set of Kraus operators $\{M_{\alpha}\}=\{\bra{\alpha}\exp(-i\delta tH_{{DS}})\ket{\phi}\}$ describes every possible indirect measurements on $S$ given the interaction (1), the state |$\ket{\phi}$ and the measurement basis {|$\{\ket{\alpha}\}$. By construction, the Kraus operators satisfy $\sum_{\alpha}M_{\alpha}^{\dagger}M_{\alpha}=I_{S}$.
@@
    A post-selection of the states is performed. This step is required depending on the chosen measurement basis, as the back action of the measured ancilla can induce an effective discontinuous evolution on |$\ket{\psi}$; that is, under the weak-measurement condition, upon expanding $M_{\alpha}$ as a power series in $\delta t$, depending on the post-selected state |$\ket{\alpha}\otimes\ket{\psi_{\alpha}}$, the latter does not evolve continuously from the initial state |$\ket{\phi}\otimes\ket{\psi}$. Therefore, the term weak comes from the (possibly) post-selected trajectories of the effectively evolved state that evolved continuously and non-unitarily from the prior state |$\ket{\psi}$.
@@ II.2 Meausurement operators
Let $\sigma,\sigma^{\prime}\in\{1,-1\},$ $\bm{\alpha}\in\{\bm{a},\bm{a}^{\prime}\}$ and $\bm{\beta}\in\{\bm{b},\bm{b}^{\prime}\},$ where $\bm{a},\bm{a}^{\prime},\bm{b},\bm{b}^{\prime}$ are fixed Bloch vectors parameterized in spherical coordinates. We define the generalized measurements acting on 4-by-4 density matrices via the following measurement operators [42, 43]
@@
where $P(\bm{\alpha},\bm{\beta})$ is the probability obtaining $(\bm{\alpha},\bm{\beta})$, and
$M_{\pm 1|\bm{\alpha}}(y)=a(\pm y)\Pi_{+1|\bm{\alpha}}+a(\mp y)\Pi_{-1|\bm{\alpha}}$ (4)
are also measurement operators. Here,
@@
is the projector associated with the Bloch vector $\bm{\alpha}=(\sin\theta_{\alpha}\cos\phi_{\alpha},\sin\theta_{\alpha}\sin\phi_{\alpha},\cos\theta_{\alpha})$, and $\bm{\sigma}$ is the vector of Pauli matrices. Furthermore,
@@
where $y$ characterizes the strength of the measurement. We only consider $y\in[0,\infty)$ for simplicity. For example, if $0<y\ll 1$, the measurement is denoted as weak. Whereas if $y=0$, then $M_{\sigma|\bm{\alpha}}(0)=I_{2}/\sqrt{2}$, where $I_{2}$ denotes the 2-by-2 identity, and it means that no measurement is performed on $\rho$. In contrast, if $y\rightarrow\infty$, i.e., (4) becomes a single projector, i.e., we obtain a strong (projective) measurement. All the above definitions also hold for $\sigma^{\prime}$ and $\bm{\beta}$.
For completeness, it is straightforward to verify that the measurement operators in (3) are well-defined for any $y$ and $z$: First, they resolve the identity in $\mathbb{C}^{4}$
@@
for an arbitrary density matrix $\rho$. Also, since we defined (6) as a positive linear combination of projectors, it implies that the measurement operators are completely positive [43], i.e., $I_{4}\otimes K^{\dagger}(\bm{\alpha},\bm{\beta};\sigma,\sigma^{\prime};y,z)K(\bm{\alpha},\bm{\beta};\sigma,\sigma^{\prime};y,z)>0$.
@@ III Bayesian Game Theory
The game-theoretic focus of this work is on Bayesian (or incomplete information) games. These were formulated classically by Harsanyi [44] and have been extended in recent years to include quantum mechanics [45, 46, 47, 48]. The essence is that players may be playing various games but are uncertain (i.e., they can have differing priors) about which game they are playing. Therefore, the probability distributions determining how the players are correlated with each other are conditioned upon which game they are playing and their subsequent choices. The conditional probability and priors determine the expected payoff for the players, where the actual payoffs can be arbitrary real numbers but are predetermined and known to both players. Importantly, there is a preference relation for these payoffs for the players associated with the game that they are playing. Throughout this paper, it is assumed that the payoff with the largest number is preferred by the players; therefore, as an example, assuming payoffs given by three and two, three is preferred over two. Subsequently, each player wishes to maximize their respective expected payoff.
In this work, we consider an incomplete information game of two players, $A$ and $B$, who share a quantum state $\rho$ describing a two-qubit system, where each player can perform local measurements represented by (4) on their subsystem. The parameters associated with players $A$ and $B$ are $\{\sigma,\bm{\alpha},y\}$ and $\{\sigma^{\prime},\bm{\beta},z\}$, respectively, where $\bm{a},\bm{a}^{\prime},\bm{b},\bm{b}^{\prime}$ are fixed measurement directions, and the strengths of the measurements $y$ and $z$ are fixed throughout the game. The incompleteness of this game is reflected in the fact that the players do not know beforehand in which direction they will be told to measure, e.g., in one trial, player $A$ can be told to measure $\bm{\alpha}=\bm{a}$ whereas player $B$ measures $\bm{\beta}=\bm{b}^{\prime}$. It is an independent referee who tells each player what direction to measure. After the measurements have been performed, the results of the measurements are communicated back to the referee, and respective payoffs are assigned. An expected (or average) payoff can be calculated for each player once this procedure has been repeated many times.
@@
where |$\ket{x}=\cos(x/2)\ket{0}+\sin(x/2)\ket{1}$ and |$\ket{0},\ket{1}$ are in the computational basis such that $\sigma^{z}\ket{0}=\ket{0}$. Other states are treated in Appendices A and B. Given the above ingredients, the expected payoff [44] in our incomplete information game is given by
@@
where $u_{A,\sigma,\sigma^{\prime}}^{\bm{\alpha},\bm{\beta}}$ denotes a tensor of payoffs for player $A$, $P(\bm{\alpha},\bm{\beta})$ denotes the prior probability, and the $\theta$’s denote the angles parameterizing the Bloch vectors associated to the measurement directions. Similar expressions hold for player $B$.
It is important to note that in our game setup, the correlation present in the state (10) can be reproduced through shared randomness by a suitable choice of a different classical correlation, as (10) is a mixed, separable state. However, in the context of measurement theory, the state is quantum, so it is ultimately affected by differing measurements, which motivates the need for the following analysis.
Table 1: Bayesian version of the Prisoner’s dilemma. Q represents quiet, and T represents tell. The police officer’s motives and the prisoner’s choices can be mapped onto ${\bm{a}},{\bm{a}^{\prime}},{\bm{b}},{\bm{b}^{\prime}},+1$ and $-1$ directly in the measurement operators (4), e.g., for prisoner $A$, Honest $\mapsto{\bm{a}}$, Corrupt $\mapsto{\bm{a^{\prime}}}$, Q $\mapsto+1$, and T $\mapsto-1$. For prisoner $B$, Honest $\mapsto{\bm{b}}$, Corrupt $\mapsto{\bm{b}^{\prime}}$, Q $\mapsto+1$, and T $\mapsto-1$. Here, prisoner $A$ chooses the rows and left entries, and prisoner $B$ chooses the columns and right entries. For example, consider that prisoner $A$ believes that the police officer is honest and measures $T$, and prisoner $B$ believes that the police officer is corrupt and measures $Q$, then prisoner $A$ gets a payoff of 3, and prisoner $B$ acquires a payoff of 2.
@@ III.1 Measurement-induced asymmetry
To clarify what asymmetry means in this context, it is instructive to understand that, by using projective measurements, the expected payoff in Bayesian game theory usually depends on both players’ measurements. Therefore, if a local measurement can be chosen such that player $A$’s expected payoff becomes independent of player $B$’s measurement, it is unanticipated that player $B$’s expected payoff still depends on player $B$’s measurements. This difference in measurement dependence on the expected payoffs is defined as asymmetry moving forward throughout this paper. This applies if players $A$ and $B$ are interchanged.
Collapsing the expected payoff so that it is independent of the other player’s measurements could have practical benefits for targeting a given value. This can be illustrated by using the Prisoner’s dilemma. The standard version of the game considers two prisoners who have been arrested for two offenses; one is a major offense (e.g., bank robbery), and the other is a minor offense (e.g., petty theft), but there is only enough evidence to charge them for the minor offense. The police would also like to charge them for the major offense, but there is insufficient evidence of this. Therefore, the police give the prisoners a choice where each prisoner can give evidence incriminating the other prisoner in exchange for a better prison sentence. However, if they both incriminate each other, then each prisoner will serve a longer prison sentence. Crucially, neither prisoner knows what the other will do.
In what follows, we consider a Bayesian version of the Prisoner’s dilemma with the added caveat that the prisoners do not know whether their respective police officer interviewing them is either honest (H) or corrupt (C).
Since the corrupt police officer does not want to get caught, the officer rewards the prisoners for confessing and taking responsibility for the crime, i.e., tell (T), which means that C will not get caught. The reward can be considered a financial payment, but the payoffs are generally arbitrary. However, if the prisoners remain quiet (Q), this increases the chance of C being caught, so the prisoners are not rewarded. A version of this type of game can be seen in Table 1.
@@
It can be seen that there is an asymmetry in the expected payoffs that arises due to the measurement when considering only one of the player’s payoffs. For example, considering prisoner $A$’s payoff. This prisoner could set $y=0$ (no measurement), which would collapse all the terms except $3/2$ in (12). Prisoner $B$ does not have this level of control over prisoner $A$’s payoff, as if $z=0$, there remain other terms that Prisoner $A$ controls. Interestingly, this asymmetry in the expected payoffs is reversed when considering the other player’s payoff in (13). Namely, prisoner $B$ does not have the same control over prisoner $A$’s payoff as prisoner $A$ if the former decides not to measure.
It is clear that in the projective limit for both players, the payoffs are symmetric in the sense that $u_{A}$ becomes $u_{B}$ when changing player $A$’s measurements to player $B$’s measurements, such that $\theta_{a}\rightarrow\theta_{b}$ and so on. This is not necessarily the case when weak measurements are considered. Despite the asymmetry in the expected payoffs due to differing types of measurement, there is still the possibility of a given prisoner optimizing the respective payoff for a given choice of measurement. For example, prisoner $A$ can choose to make no measurement ($y=0$) and fix the expected payoff at $3/2$. Despite this, prisoner $B$ would still be left to control the measurements for the payoff $u_{B}$. Thus, in theory, prisoner B would be able to acquire a larger expected payoff than prisoner $A$ by choosing the detector settings ($\theta_{b},\theta_{b^{\prime}}$) in an optimal way and vice versa for prisoner $B$. Subsequently, each player (if they so choose) has complete control over their respective expected payoff and can make their expected payoff completely independent of the other player’s choices (measurements) while allowing the other player to still have control over the corresponding particular payoff. This could have an application where one player is more risk-averse compared to the other, so this player can choose whether to allow the other player to influence their respective payoff.